# TLDP: An Algorithm of Local Differential Privacy for Tensors

### v1

@@ V-B1 Type I: Private Testing Data
The datasets used in Type I include MNIST, CIFAR-10, SVHN, and IMDB. The setup is in Table II. In this type, every segment of the training data is private, crucial, and must be preserved entirety. Users manipulate their respective data and send it directly to a central server, which in turn uses this information to train the model parameters before transmitting the updated parameters back. No preprocessing or size-trimming is applied to the raw data, ensuring that the accuracy of the experimental results is maintained at the highest possible level.
Additionally, for the MNIST dataset, we include comparisons with the TLDP-L-w and TLDP-G-w methods, where TLDP-L-w and TLDP-G-w represent the proposed TLDP when using Laplace or Gaussian noise while applying a weight matrix for improving utility. In these approaches, the weights are determined based on the grayscale values of the images: higher grayscale values correspond to higher weights. This setting is reasonable since, as the dataset has been preprocessed, apart from the digits that have values greater than 0, all other regions are set to 0.
Query function: The query function we use is the identity function $f(\mathcal{X})=\mathcal{X}$, where $\mathcal{X}$ represents the original tensors from clients. Since TVG is a DP mechanism, we need to define $l_{2}$ sensitivity first, for neighboring datasets $\{\mathcal{X},\mathcal{X}^{\prime}\}$:
$s_{2}(f)=\underset{\mathcal{X},\mathcal{X}^{\prime}}{sup}||\mathcal{X}-\mathcal{X}^{\prime}||_{F}=2\sqrt{I_{1}I_{2}\dots I_{N}},$
@@
TABLE II: SETUP FOR TESTING DATA
Dataset MNIST Cifar-10 SVHN IMDB
Model MobileNet MobileNet MobileNet BiLSTM
Training Data 60000 50000 73257 25000
Testing Data 10000 10000 26032 25000
Batch size 64 64 64 64
V-B2 Type II: Private Training Features
For Type II, we select MNIST, CIFAR-10, and SVHN as experimental datasets. The setup is shown in Table III. DarkNet and IMDB are not used in this task due to the absence of pre-trained models. We consider all training data for these datasets to be confidential and are committed to ensuring privacy. Given that high-dimensional data can amplify the noise added during the process, we instead transmit low-dimensional features extracted by shallow neural networks. To enhance model performance, we substitute the activation function in each layer of MobileNet from ReLU6 to tanh()($\cdot$)( ⋅ ), ensuring that the output features are normalized within the range of $(-1,1)$. MobileNet is divided into two parts: the feature extractor (comprising all layers except the final fully connected layer) and the classifier (the final fully connected layer). After extracting features from the feature extractor, noise is added to the features, and only the classifier is trained and tested.
Query function: We use the identity function $f(\mathcal{X})=\mathcal{X}$ as the query function, and $\mathcal{X}$ represent features extraced from the original data. In TVG, for neighboring datasets $\{\mathcal{X},\mathcal{X}^{\prime}\}$, the $l_{2}$ sensitivity is:
$s_{2}(f)=\underset{\mathcal{X},\mathcal{X}^{\prime}}{sup}||\mathcal{X}-\mathcal{X}^{\prime}||_{F}=2\sqrt{I_{1}I_{2}\dots I_{N}},$
since the range of data is set to $(-1,1)$.
TABLE III: Setup for Type II: private training feature
@@ V-B3 Type III: Private SGD
Datasets include MNIST, CIFAR-10, SVHN, IMDB, and DarkNet, as shown in Table IV. In the experiments, there are three local models and one global model. The training data is split into three parts, with each local model training on one subset. After completing a single epoch, each local model uploads its weights and biases to the global model. The global model then updates its parameters by averaging the weights and biases from the three local models and subsequently sends the updated parameters back to the local models. As a result, after each iteration, the weights and biases of all four models are synchronized. We employ the noisy SGD algorithm as the local differentially private optimization method. The entire process can be restated as follows:
  • Compute the gradients for this subset.
  • Clip the gradients by their $L_{\infty}$ norm, with a clipping threshold of $C$.
  • Calculate the mean of the gradients across a batch of subsets and introduce noise to these averaged gradients to ensure privacy.
  • Update the model parameters using these noise-adjusted gradients and repeat the process starting from step 1.
Query function: The query function we use is the identity function $f(\mathcal{X})=\mathcal{X}$, and $\mathcal{X}$ represents the gradients trained by clients. Similarly, we define $l_{2}$ sensitivity as, for neighboring datasets $\{\mathcal{X},\mathcal{X}^{\prime}\}:$
$s_{2}(f)=\underset{\mathcal{X},\mathcal{X}^{\prime}}{sup}||\mathcal{X}-\mathcal{X}^{\prime}||_{F}=2C\sqrt{I_{1}I_{2}\dots I_{N}}$
@@
TABLE IV: SETUP FOR Private SGD)
@@ V-C Experimental Results
Method Laplace Gaussian RR Laplace RR Gaussian IDN
$\mathbb{E}||\mathcal{Z}||_{2}$ $\frac{\Delta}{\epsilon}I\sqrt{2I}$ $\frac{\Delta}{2\epsilon}I\sqrt{I}$ $\frac{\Delta}{\epsilon}\sqrt{2(1-p)I}$ $\frac{\Delta}{2\epsilon}\sqrt{(1-p)I}$ $\sqrt{\frac{Is_{2}^{2}(f)}{(-\zeta(\delta)+\sqrt{\zeta^{2}(\delta)+2\epsilon})^{2}}}$
Type I Results: Fig. 1 summarizes the results of Type I across various datasets when $\epsilon=10$. It demonstrates that in computer vision tasks, our TLDP algorithm is able to achieve high classification accuracy, while other methods struggle due to the excessive noise introduced by the large scale of the data, making accurate identification nearly impossible.
In this table, $I=I_{1}I_{2}\dots I_{N}$ is the product of each dimension of the tensor. Since the query functions for the three tasks outlined in Section V-B are identical, the expected errors are consistent, which is defined as $\underset{\mathcal{Z}\in\mathbb{R}^{I_{1}I_{2}\dots I_{N}}}{sup}\mathbb{E}||\mathcal{Z}||_{2}=\underset{\mathcal{Z}\in\mathbb{R}^{I_{1}I_{2}\dots I_{N}}}{sup}\mathbb{E}(\sum_{i=1}^{I}(\mathcal{Z}_{i})^{2})^{\frac{1}{2}}$, where $\mathcal{Z}_{i}(i=1,2,\dots,N)$ denote the components of $\mathcal{Z}$. Using the properties of variance, we can derive that for TLDP-L noise (TLDP-L denotes TLDP that uses the Laplace noise),
@@
$\underset{\mathcal{Z}\in\mathbb{R}^{I_{1}I_{2}\dots I_{N}}}{sup}\mathbb{E}||\mathcal{Z}||_{2}=\frac{\Delta}{\epsilon}\sqrt{2(1-p)I}.$
For TLDP-G noise (TLDP-G represents TLDP that uses Gaussian noise),
@@
$\underset{\mathcal{Z}\in\mathbb{R}^{I_{1}I_{2}\dots I_{N}}}{sup}\mathbb{E}||\mathcal{Z}||_{2}=\frac{\Delta}{2\epsilon}\sqrt{(1-p)I}.$
If the randomized response mechanism is removed from our method, $i.e.$ no component will be retained with the probability $p$, and the term $1-p$ in the outcomes will become $1$. Moreover, as discussed in the proof of Theorem 2, in order to ensure the same privacy budget, the values of the two noise parameters are increased by a factor of I compared to the improved methods. As a result, the expected errors of normal Laplace and Gaussian mechanisms are $\frac{\Delta}{\epsilon}I\sqrt{2I}$ and $\frac{\Delta}{2\epsilon}I\sqrt{I}$ respectively.
For independent noise (IDN), we adopt the same parameters as Task II in [11], which provides the expected error $\frac{Is_{2}^{2}(f)}{(-\zeta(\delta)+\sqrt{\zeta^{2}(\delta)+2\epsilon})^{2}}$, where $\zeta^{2}(\delta)=-2\ln{\delta}+2\sqrt{-I\ln{\delta}}+I$. Removing the relaxation term $\delta$, we obtain its estimated value $\frac{\Delta I\sqrt{I}}{\sqrt{2}\epsilon}.$
@@
Fig. 1 reports F1-score under different $\epsilon$, and the relaxation term $\delta$ is fixed as $10^{-5}$, similar to [11]. The experimental results across three datasets clearly demonstrate that as $\epsilon$ increases, the F1-scores of the Laplace, Gaussian, MVG, and IDN mechanisms remain consistently low, indicating their inability to effectively complete the task. In contrast, the F1-score of the TLDP method rises sharply. On the MNIST and SVHN datasets, the score reaches approximately 90% when $\epsilon$ is below 1. Although the growth is slower on the CIFAR-10 dataset, it eventually stabilizes around 70%. For the MNIST dataset, we also included the TLDP-L-w and TLDP-G-w methods. While the performance of the former is nearly identical to that of TLDP-L, the latter slightly underperforms TLDP-G, which is consistent with our earlier theoretical analysis. The addition of a weight matrix ensures better privacy protection under the same differential privacy budget, at the cost of slightly reduced accuracy. Furthermore, the TLDP-G mechanism outperforms the TLDP-L mechanism, in line with our previous theoretical error analysis.
Type II Results: The results of Type II for each dataset are presented in Fig. 2 when $\epsilon=10$. Due to the fact that the central server now operates on extracted feature tensors instead of raw data in the training set, the performance metrics of the TLDP method experience a decline. However, despite this reduction, the TLDP method still significantly outperforms the Laplace, Gaussian, MVG, and IDN mechanisms.
@@
Type III Results: Fig. 3 shows the results of Type III on the different datasets for comparison when $\epsilon=140$. Two new datasets, IMDB for data mining and DarkNet for text mining, are introduced in the experiment. In Type III, due to the transmission of more abstract gradients to the central server, the results from the various methods only become comparable when $\epsilon$ exceeds 100. As anticipated, the performance of the Laplace, Gaussian, MVG, and IDN mechanisms remains significantly low. In contrast, the TLDP method consistently delivers superior results, particularly on the MNIST, SVHN, IMDB, and DarkNet datasets.
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (Grant No. 62406215) and the ”Fundamental Research Funds for the Central Universities”, This work was also supported in part by the National Natural Science Foundation of China (62072321), the Six Talent Peak Project of Jiangsu Province (XYDXX-084), the Science and Technology Program of Jiangsu Province (BZ2024062), the Natural Science Foundation of the Jiangsu Higher Education Institutions of China (22KJA520007), Suzhou Planning Project of Science and Technology (2023ss03).

### v2

@@ V-B1 Type I: Private Testing Data
The datasets used in Type I include MNIST, CIFAR-10, and SVHN. The setup is illustrated in Table II. In this type, the training data is likely to be private and should be preserved. Users manipulate their respective data and send it directly to a central server, which uses this information to train the model parameters before transmitting the updated parameters back. No preprocessing or size-trimming is applied to the raw data, ensuring that the accuracy of the experimental results is maintained at the highest possible level.
Additionally, for the MNIST dataset, we include comparisons with the TLDP-L-w and TLDP-G-w methods. In these approaches, the weights are determined based on the grayscale values of the images: higher grayscale values correspond to higher weights. This setting is reasonable since, as the dataset has been preprocessed, apart from the digits with values greater than 0, all other regions are set to 0.
Query function: The query function we use is the identity function $f(\mathcal{X})=\mathcal{X}$, where $\mathcal{X}$ represents the original tensors from users. The $l_{2}$ sensitivity for neighboring datasets $\{\mathcal{X},\mathcal{X}^{\prime}\}$ is defined as:
$s_{2}(f)=\underset{\mathcal{X},\mathcal{X}^{\prime}}{sup}||\mathcal{X}-\mathcal{X}^{\prime}||_{F}=2\sqrt{I_{1}I_{2}\dots I_{N}},$ (17)
@@
TABLE II: Datasets and parameter settings for Type I
@@
$\Delta$ 256 256 256
V-B2 Type II: Private Training Features
For Type II, we select MNIST, CIFAR-10, and SVHN as experimental datasets. The setup is shown in Table III. DarkNet and IMDB are not used in this task due to the absence of pre-trained models, and without them, the extracted feature embeddings are prone to be less meaningful. We consider the features extracted from the original training data of these datasets to be confidential and are committed to protecting privacy. To enhance model performance, we substitute the activation function in each layer of MobileNet from ReLU6 to tanh()($\cdot$)( ⋅ ), ensuring that the output features are normalized within the range of $(-1,1)$. MobileNet is divided into two parts: the feature extractor (i.e., all layers except the final fully connected layer) and the classifier (i.e., the final fully connected layer). After extracting features from the feature extractor, noise is added to the features, and only the classifier is trained and tested.
Query function: We use the identity function $f(\mathcal{X})=\mathcal{X}$ as the query function, and $\mathcal{X}$ represent features extracted from the original data. In TVG, for neighboring datasets $\{\mathcal{X},\mathcal{X}^{\prime}\}$, the $l_{2}$ sensitivity is:
$s_{2}(f)=\underset{\mathcal{X},\mathcal{X}^{\prime}}{sup}||\mathcal{X}-\mathcal{X}^{\prime}||_{F}=2\sqrt{I_{1}I_{2}\dots I_{N}},$ (18)
since the range of data is set to $(-1,1)$.
TABLE III: Datasets and parameter settings for Type II
Dataset MNIST Cifar-10 SVHN
Model MobileNet MobileNet MobileNet
Training Data 60000 50000 73257
Testing Data 10000 10000 26032
Batch size 64 64 64
$\Delta$ 256 256 256
@@ V-B3 Type III: Private SGD
For this task, we utilized datasets including MNIST, CIFAR-10, SVHN, IMDB, and DarkNet, as shown in Table IV. The training data is split into three parts, and each local model is trained on one subset. After completing a single epoch, each local model uploads its weights and biases using the global model. The global model then updates its parameters by averaging the weights and biases from the three local models and subsequently sends the updated parameters back to the local models. As a result, after each iteration, the weights and biases of all four models are synchronized. The entire process can be restated as follows:
  • Train each local model on its subset.
  • Clip the model parameters by their $L_{\infty}$ norm, with a clipping threshold of $C$.
  • Calculate the mean of the model parameters across a batch of subsets and introduce noise to these averaged model parameters to ensure privacy.
  • Update the model parameters using these noise-adjusted model parameters and repeat the process starting from step 1.
Query function: The query function we use is the identity function $f(\mathcal{X})=\mathcal{X}$, and $\mathcal{X}$ represents the model parameters trained by users. Similarly, we define $l_{2}$ sensitivity as, for neighboring datasets $\{\mathcal{X},\mathcal{X}^{\prime}\}:$
$s_{2}(f)=\underset{\mathcal{X},\mathcal{X}^{\prime}}{sup}||\mathcal{X}-\mathcal{X}^{\prime}||_{F}=2C\sqrt{I_{1}I_{2}\dots I_{N}}$ (19)
@@
TABLE IV: Datasets and parameter settings for Type III
@@ V-C Experimental Results
Method Laplace Gaussian MVG IDN TLDP-L TLDP-G
$\mathbb{E}||\mathcal{Z}||_{2}$ $\frac{\Delta}{\epsilon}I\sqrt{2I}$ $\frac{\Delta}{2\epsilon}I\sqrt{I}$ $\frac{I_{1}I}{(-\beta_{0}+\sqrt{\beta_{0}^{2}+8\alpha_{0}\epsilon})^{2}}\cdot 4\alpha_{0}^{2}$ $\sqrt{\frac{Is_{2}^{2}(f)}{(-\zeta(\delta)+\sqrt{\zeta^{2}(\delta)+2\epsilon})^{2}}}$ $\frac{\Delta}{\epsilon}\sqrt{2(1-p)I}$ $\frac{\Delta}{2\epsilon}\sqrt{(1-p)I}$
In Table V, $I=I_{1}I_{2}\dots I_{N}$ is a product of each dimension of the tensor. Since the query functions for the three tasks outlined in Section V-B are identical, the expected errors are consistent, which is defined as
$\underset{\mathcal{Z}\in\mathbb{R}^{I_{1}I_{2}\dots I_{N}}}{sup}\mathbb{E}||\mathcal{Z}||_{2}=\underset{\mathcal{Z}\in\mathbb{R}^{I_{1}I_{2}\dots I_{N}}}{sup}\mathbb{E}(\sum_{i=1}^{I}(\mathcal{Z}_{i})^{2})^{\frac{1}{2}},$ (20)
where $\mathcal{Z}_{i}(i=1,2,\dots,N)$ denote the components of $\mathcal{Z}$. Using the properties of variance, we can derive that for TLDP-L noise, we have
${}^{2}\leq\mathbb{E}[(\sum_{i=1}^{I}(\mathcal{Z}_{i})^{2})^{\frac{1}{2}}]^{2}=\mathbb{E}(\sum_{i=1}^{n}(\mathcal{Z}_{i})^{2})$ (21)
@@
$\underset{\mathcal{Z}\in\mathbb{R}^{I_{1}I_{2}\dots I_{N}}}{sup}\mathbb{E}||\mathcal{Z}||_{2}=\frac{\Delta}{\epsilon}\sqrt{2(1-p)I}.$ (22)
For TLDP-G noise,
@@
$\underset{\mathcal{Z}\in\mathbb{R}^{I_{1}I_{2}\dots I_{N}}}{sup}\mathbb{E}||\mathcal{Z}||_{2}=\frac{\Delta}{2\epsilon}\sqrt{(1-p)I}.$ (23)
If the randomized response mechanism is removed from our method, i.e., no component will be retained with the probability $p$, and the term $1-p$ in the outcomes will become $1$. Moreover, as discussed in the proof of Theorem 2, in order to ensure the same privacy budget, the values of the two noise parameters are increased by a factor of $I$ compared to the improved methods. As a result, the expected errors of normal Laplace and Gaussian mechanisms are $\frac{\Delta}{\epsilon}I\sqrt{2I}$ and $\frac{\Delta}{2\epsilon}I\sqrt{I}$, respectively.
For MVG and IDN, we adopt the same parameters as Task II in [16] and [17], which provides the expected error $\frac{I_{1}I}{(-\beta_{0}+\sqrt{\beta_{0}^{2}+8\alpha_{0}\epsilon})^{2}}\cdot 4\alpha_{0}^{2}$ and $\sqrt{\frac{Is_{2}^{2}(f)}{(-\zeta(\delta)+\sqrt{\zeta^{2}(\delta)+2\epsilon})^{2}}}$ (where $\zeta^{2}(\delta)=-2\ln{\delta}+2\sqrt{-I\ln{\delta}}+I$). Removing the relaxation term $\delta$, we obtain its estimated value $\frac{\Delta I_{1}I\sqrt{I}}{\sqrt{2}\epsilon}\cdot\ln^{2}(I_{1}+1)$ and $\frac{\Delta I\sqrt{I}}{\sqrt{2}\epsilon}.$
@@
V-C1 Type I Results
Fig. 1 summarizes the results of Type I across various computer vision datasets, i.e., MNIST, Cifar-10, and SVHN. It demonstrates that in computer vision tasks, our TLDP algorithm can achieve much higher classification accuracy than the comparative methods because the latter introduces excessive noise to the tensor data, making accurate identification nearly impossible.
Fig. 1 reports F1-score under different $\epsilon$, and the relaxation term $\delta$ is fixed as $10^{-5}$, similar to [17]. The experimental results across three datasets clearly demonstrate that as $\epsilon$ increases, the F1-scores of the Laplace, Gaussian, MVG, and IDN mechanisms remain consistently low, indicating their inability to effectively complete the task. In contrast, the F1-score of the TLDP method rises sharply. On the MNIST and SVHN datasets, the score reaches approximately 90% when $\epsilon$ is below 1. Although the growth is slower on the CIFAR-10 dataset, it eventually stabilizes around 70%. For the MNIST dataset, we also included the TLDP-L-w and TLDP-G-w methods. While the performance of the former is nearly identical to that of TLDP-L, the latter slightly underperforms TLDP-G, which is consistent with our earlier theoretical analysis. The addition of a weight matrix ensures better privacy protection under the same differential privacy budget at the cost of slightly reduced accuracy. Furthermore, the TLDP-G mechanism outperforms the TLDP-L mechanism, which is in line with our previous theoretical error analysis.
V-C2 Type II Results
The results of Type II for each dataset are presented in Fig. 2. Due to the fact that the central server now operates on extracted feature tensors instead of raw data in the training set, the performance metrics of the TLDP method experience a decline. However, despite this reduction, the TLDP method still significantly outperforms the Laplace, Gaussian, MVG, and IDN mechanisms.
@@ V-C2 Type II Results
V-C3 Type III Results
Fig. 3 shows the results of Type III on the different datasets for comparison. Two new datasets, IMDB for text mining and DarkNet for data mining, are introduced in the experiment. In Type III, due to the transmission of more abstract model parameters to the central server, the results from the various methods only become comparable when $\epsilon$ exceeds 100. As anticipated, the performance of the Laplace, Gaussian, MVG, and IDN mechanisms remains significantly low. In contrast, the TLDP method consistently delivers superior results, particularly on the MNIST, SVHN, IMDB, and DarkNet datasets.
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (Grant No. 62406215 and 62072321).